-LEAVE ANSWERS OUT TO 4 DECIMAL PLACES

- PLEASE SHOW WORK

-EACH PART OF EACH QUESTION IS WORTH 2 POINTS FOR A TOTAL OF 38 POINTS.

1. Bank account balance is an example of a variable that follows which scale of measurement?

            a. ratio scale

            b. interval scale

            c. nominal scale

            d. ordinal scale

2./The number of lighting deaths per state in the Midwest over 20 years is listed in the table below. We are interested in fully analyzing the number of lighting deaths.

State

# of Lighting Deaths over 20 years

Illinois

11

Indiana

6

Iowa

3

Kansas

2

Michigan

10

Minnesota

2

Missouri

7

Nebraska

4

North Dakota

1

Ohio

13

South Dakota

3

Wisconsin

8

a. Is this data categorical or quantitative? If quantitative, is it discrete or continuous?

            b. Calculate the number of classes and class width.

            c. Construct a frequency distribution, including the classes, frequency, and relative frequency.

            d. Construct a histogram or bar graph (depending on your answer to part a), including all labels.

            e. Calculate the mean, median, and mode.

            f. Are the mean, median, and mode descriptive or inferential statistics?

            g. Calculate the range.

h. Calculate Q1.

i. Calculate Q3.

j. Calculate the IQR.

k. Find the variance.

            l. Find the standard deviation.

            m. Find the coefficient of variation.

            n. What is the 80th percentile of this data set?

            o. Does this data set have any outliers? Use statistics in answering this question (not just your opinion).

            p. Would you use the Empirical rule or Chebyshev’s Theorem here and why?

            q. Based on the mean and median calculated above, would you expect this data to be symmetric, skewed to the left or skewed to the right and why?

           3. A survey of 500 people in Chicago was taken to compare how many are Cubs vs. Sox fans by gender. The cross tabulation is listed below:

Gender

Cubs Fans

Sox Fans

Total

Male

250

90

340

Female

100

60

160

Total

350

150

500

Of the people surveyed who are male, what percentage are Cubs fans?
Answer #1

Solution1:

Bank account balance is an example of ratio scale of measurement ,in addition to observations having equal intervals,can have a value of zero as well.

2./The number of lighting deaths per state in the Midwest over 20 years is listed in the table below. We are interested in fully analyzing the number of lighting deaths.

we are measuring data it is quantitative. Data does not have integers it is discrete

a. Is this data categorical or quantitative? If quantitative, is it discrete or continuous?

QUANTITATIVE,DISCRETE

b. Calculate the number of classes and class width.

Number of classes=k=1+3.322logn

=1+3.322log12

=4.585

=5

number of classes=5(rounding to next integer , no of classes cannot have decimals)

class width=Range/no of classes

=13-1/5

= 12/5

=2.4

class width=3

Class interval Frequency Relative frequency=frequenc/total
1-3.4 5 0.416666667
3.4-5.8 1 0.083333333
5.8-8.2 3 0.25
8.2-10.6 1 0.083333333
10.6-13 2 0.166666667
Total' 12 1

Solution-4:

in excel select the data

go to insert>histogram

Histogram [1, 3.4] (3.4,5.8] (5.8, 8.2] (8.2, 10.6) (10.6, 13]